OpenStudy (anonymous):

Is the right end of equation 9.1.5 correct? If I do the integration I get cos(theta2) - cos(theta1) because cos(-theta1) = cos(theta1)

OpenStudy (anonymous):

Ok I think I've figured it out after looking at it for an hour or 2, the integral should be \[\frac{ -\mu _{0} . I}{ 4\pi a } \int\limits_{\pi - \theta_{1}}^{\theta _{2}}\sin \theta \delta \Theta \] to get correct signs. It also seems more logical when working with these angels